# Let’s Be Self-generated via Step by Step: A Curriculum Learning Approach to Automated Reasoning with Large Language Models

### v1

@@ 5 Conclusion
Discussions. Here, we discuss the shortcomings of the LBS3 method as follows:
@@
2) The proposed LBS3 suggests that LLMs generate both easy and hard queries, but it does not delve into a clear definition of whether the generated queries are genuinely easy or hard. Existing works (Yasunaga et al., 2023; Yang et al., 2023; Chen et al., 2023) similarly lack research in this area. We believe that the aforementioned analysis is necessary to ascertain whether the reported improvements are truly because the queries have become easier (for humans or models) and more hard, or simply due to the prompts. To this end, in Appendix C, we provide examples of simple and complex proxy queries generated by five LLMs, as shown in Tables 2 to 49. It can be intuitively observed that they can generate simple and difficult queries based on prompts, thereby intuitively confirming the main claim of this paper related to curriculum learning. However, the black-box nature of Large Language Models (LLMs) precludes us from conducting a comprehensive qualitative analysis, even though we are keen to do so. In summary, the effectiveness of our method—and related approaches, including Analogical Prompting, Self-ICL, and Auto-ICL—is predicated on the LLM’s capacity to follow instructions and its a wealth of knowledge that enable them to fulfill various reasoning tasks. We will continue to address the aforementioned shortcomings in future work.
Broader Impacts. We focus on how to guide LLMs to effectively solve a given query in scenarios without any additional information. Our work reveals that existing approaches for such scenarios either require solving multiple proxy queries of similar or even greater difficulty, leading to mediocre proxy exemplars prompts, or place high demands on the LLMs’ ability to follow instructions and generate responses. Our proposed LBS3 approch successfully alleviates the above issues. LBS3 embodies potential positive social impacts by realizing a prompting framework with exceptional performance, offering insights for real-world prompt engineering applications. Also, LBS3 may have negative social impacts related to sensitive information and high resource consumption. In addition, LBS3 approach, based on open-source LLMs, requires significant electrical resources for executing reasoning tasks in bulk. LBS3 does not involve social ethics.
@@
Appendix A Related Works
@@
Curriculum Learning. The underlying insight of curriculum learning is to emulate the learning paradigms of humans and animals, that is, by following the sequence and content of standardized educational materials, they leverage previously learned concepts to aid in the acquisition of new and more challenging ones (Krueger & Dayan, 2009; Pavlov, 2010; Skinner, 1958). Inspired by cognitive science research (Rohde & Plaut, 1999), curriculum-based machine learning algorithm was first proposed by (Bengio et al., 2009) with the core idea of initially training models with simple samples and gradually increasing the complexity during the training process. Over the subsequent decade, the concept of curriculum learning has been widely applied in the field of artificial intelligence, including computer vision (Gong et al., 2016; Xiangli et al., 2022; Yang et al., 2022; Zhou et al., 2022c), machine translation (Zhang et al., 2018; Platanios et al., 2019; Liu et al., 2020; Mohiuddin et al., 2022),pre-training (Campos, 2021; Li et al., 2021; Nagatsuka et al., 2021; Zhou et al., 2022a), fine-tuning (Chen et al., 2024b; Gao et al., 2024; Yigit & Amasyali, 2023; Zhang et al., 2022), natural language understanding (Wang et al., 2022a; Zhang et al., 2021; Xu et al., 2020; Christopoulou et al., 2022), knowledge distillation (Li et al., 2023b; a; Matiisen et al., 2019; Zhu et al., 2021; Maharana & Bansal, 2022), and more. However, the utilize of curriculum learning strategies to enhance the reasoning capabilities of language models remains unexplored. To the best of our knowledge, our work is a pioneering attempt to mimic the idea of curriculum learning, aiming to investigate how LLMs can self-generate few-shot exemplary prompts to facilitate the reasoning process.
@@
Chain of thought (CoT) Prompting Approaches. Few-Shot CoT (FS-CoT), initially proposed by Wei et al. (Wei et al., 2022b), has shown that providing intermediate reasoning steps (termed "thoughts") in manually crafted few-shot exemplary prompts can ignite the step-by-step reasoning capabilities of LLMs, thereby significantly enhancing their accuracy in solving complex reasoning tasks. This approach is bolstered by the self-consistency approach (Wang et al., 2022b; Aggarwal et al., 2023b; Li et al., 2024). Despite its achievements, Few-Shot CoT confronts challenges such as the accumulation of errors, the limited quality of exemplary prompts, and the time-consuming labor-intensive task of manual annotation.
@@
In order to alleviate the performance degradation caused by accumulated errors, a plethora of variants for Few-shot CoT have been proposed. For instance, there are more complex CoT approaches (Lee & Kim, 2023; Chen et al., 2024a; Yao et al., 2024; Besta et al., 2024; Zou et al., 2023; Yu et al., 2024; Zhou et al., 2022b; Sun et al., 2023; Wang et al., 2023; Yin et al., 2023a; Zhao et al., 2023) as well as those with feedback and verification mechanisms (Zhang et al., 2023; Ling et al., 2024; Poesia et al., 2023; Paul et al., 2023; Weng et al., 2022; Madaan et al., 2024), etc. The mentioned methods are committed to constructing frameworks that guide the language model to generate correct intermediate steps, thereby reducing accumulated errors in the intermediate reasoning process and improving the accuracy of the final answer. However, such meticulously designed frameworks inevitably come with a steep computational cost.
@@
Research indicates that existing LLMs are sensitive to the quality and sequence of exemplary prompts, making the construction of high-quality prompts crucial (Liu et al., 2021; Lu et al., 2021). Consequently, a series of efforts have been dedicated to enhancing the quality of these exemplary prompts (Rubin et al., 2021; Fu et al., 2022; Ye et al., 2022; Su et al., 2022; Wu et al., 2022; Ye & Durrett, 2023; Diao et al., 2023; Wan et al., 2023). The above-mentioned approaches rest on a fundamental assumption that there is an accessible external resource related to the current task, such as a dataset or corpus. They employ various predefined similarity metrics to retrieve the most relevant, complex and diverse existing queries or exemplars from the external resource to improve the quality of exemplary prompts. Nevertheless, the requisite external resources these approaches rely on are not always available in practice, and they may not entirely circumvent the need for manual annotation.
@@ Appendix B Computing Devices and Platforms
The following is the configuration of the computing device for our experiments using open-source LLMs.

### v3

@@ 5 Conclusion
Limitations
Here, we discuss the shortcomings of the LBS3 method as follows:
@@
2) The proposed LBS3 suggests that LLMs generate both easy and hard queries, but it does not delve into a clear definition of whether the generated queries are genuinely easy or hard. Existing works (Yasunaga et al., 2023; Yang et al., 2023; Chen et al., 2023) similarly lack research in this area. We believe that the aforementioned analysis is necessary to ascertain whether the reported improvements are truly because the queries have become easier (for humans or models) and more hard, or simply due to the prompts. To this end, in Appendix C, we provide examples of simple and complex proxy queries generated by five LLMs, as shown in Tables 4 to 52. It can be intuitively observed that they can generate simple and difficult queries based on prompts, thereby intuitively confirming the main claim of this paper related to curriculum learning. However, the black-box nature of Large Language Models (LLMs) precludes us from conducting a comprehensive qualitative analysis, even though we are keen to do so. In summary, the effectiveness of our method—and related approaches, including Analogical Prompting, Self-ICL, and Auto-ICL—is predicated on the LLM’s capacity to follow instructions and its a wealth of knowledge that enable them to fulfill various reasoning tasks. We will continue to address the aforementioned shortcomings in future work.
Ethical Considerations
We focus on how to guide LLMs to effectively solve a given query in scenarios without any additional information. Our work reveals that existing approaches for such scenarios either require solving multiple proxy queries of similar or even greater difficulty, leading to mediocre proxy exemplars prompts, or place high demands on the LLMs’ ability to follow instructions and generate responses. Our proposed LBS3 approch successfully alleviates the above issues. LBS3 embodies potential positive social impacts by realizing a prompting framework with exceptional performance, offering insights for real-world prompt engineering applications. Also, LBS3 may have negative social impacts related to sensitive information and high resource consumption. In addition, LBS3 approach, based on open-source LLMs, requires significant electrical resources for executing reasoning tasks in bulk. LBS3 does not involve social ethics.
@@
Appendix A Related Work
@@
Curriculum Learning. The underlying insight of curriculum learning is to emulate the learning paradigms of humans and animals, that is, by following the sequence and content of standardized educational materials, they leverage previously learned concepts to aid in the acquisition of new and more challenging ones (Krueger and Dayan, 2009; Pavlov, 2010; Skinner, 1958). Inspired by cognitive science research (Rohde and Plaut, 1999), curriculum-based machine learning algorithm was first proposed by (Bengio et al., 2009) with the core idea of initially training models with simple samples and gradually increasing the complexity during the training process. Over the subsequent decade, the concept of curriculum learning has been widely applied in the field of artificial intelligence, including computer vision (Gong et al., 2016; Xiangli et al., 2022; Yang et al., 2022; Zhou et al., 2022c), machine translation (Zhang et al., 2018; Platanios et al., 2019; Liu et al., 2020; Mohiuddin et al., 2022),pre-training (Campos, 2021; Li et al., 2021; Nagatsuka et al., 2021; Zhou et al., 2022a), fine-tuning (Chen et al., 2024b; Gao et al., 2024; Yigit and Amasyali, 2023; Zhang et al., 2022), natural language understanding (Wang et al., 2022a; Zhang et al., 2021; Xu et al., 2020; Christopoulou et al., 2022), knowledge distillation (Li et al., 2023b, a; Matiisen et al., 2019; Zhu et al., 2021; Maharana and Bansal, 2022), and more. However, the utilize of curriculum learning strategies to enhance the reasoning capabilities of language models remains unexplored. To the best of our knowledge, our work is a pioneering attempt to mimic the idea of curriculum learning, aiming to investigate how LLMs can self-generate few-shot exemplary prompts to facilitate the reasoning process.
@@
Chain of thought (CoT) Prompting Approaches. LLMs have been explored by researchers for various tasks due to their powerful emergent capabilities Yu et al. (2024a); Tan et al. (2024); Si et al. (2025); Liu et al. (2025); Luo et al. (2025). Few-Shot CoT (FS-CoT), initially proposed by Wei et al. (Wei et al., 2022b), has shown that providing intermediate reasoning steps (termed "thoughts") in manually crafted few-shot exemplary prompts can ignite the step-by-step reasoning capabilities of LLMs, thereby significantly enhancing their accuracy in solving complex reasoning tasks. This approach is bolstered by the self-consistency approach (Wang et al., 2022b; Aggarwal et al., 2023b; Li et al., 2024). Despite its achievements, Few-Shot CoT confronts challenges such as the accumulation of errors, the limited quality of exemplary prompts, and the time-consuming labor-intensive task of manual annotation.
@@
In order to alleviate the performance degradation caused by accumulated errors, a plethora of variants for Few-shot CoT have been proposed. For instance, there are more complex CoT approaches (Lee and Kim, 2023; Chen et al., 2024a; Yao et al., 2024; Besta et al., 2024; Zou et al., 2023; Yu et al., 2024b; Zhou et al., 2022b; Sun et al., 2023; Wang et al., 2023; Yin et al., 2023a; Zhao et al., 2023) as well as those with feedback and verification mechanisms (Zhang et al., 2023; Ling et al., 2024; Poesia et al., 2023; Paul et al., 2023; Weng et al., 2022; Madaan et al., 2024), etc. The mentioned methods are committed to constructing frameworks that guide the language model to generate correct intermediate steps, thereby reducing accumulated errors in the intermediate reasoning process and improving the accuracy of the final answer. However, such meticulously designed frameworks inevitably come with a steep computational cost.
@@
Research indicates that existing LLMs are sensitive to the quality and sequence of exemplary prompts, making the construction of high-quality prompts crucial (Liu et al., 2021; Lu et al., 2021). Consequently, a series of efforts have been dedicated to enhancing the quality of these exemplary prompts (Rubin et al., 2021; Fu et al., 2022; Ye et al., 2022; Su et al., 2022; Wu et al., 2022; Ye and Durrett, 2023; Diao et al., 2023; Wan et al., 2023). The above-mentioned approaches rest on a fundamental assumption that there is an accessible external resource related to the current task, such as a dataset or corpus. They employ various predefined similarity metrics to retrieve the most relevant, complex and diverse existing queries or exemplars from the external resource to improve the quality of exemplary prompts. Nevertheless, the requisite external resources these approaches rely on are not always available in practice, and they may not entirely circumvent the need for manual annotation.
@@ Appendix B Computing Devices and Platforms
The following is the configuration of the computing device for our experiments using open-source LLMs. Our code is here: https://anonymous.4open.science/r/LBS3-B926.
@@ Appendix C Examples of Proxy Queries Generated with Varying Prompt Templates
Appendix D Research on the Quality of Easy and Hard-Proxy Queries for LBS3
In this section, we first provide scoring rules of difficulty for proxy queries using GPT-4o, as shown in Table 3. Additionally, to gain a more detailed understanding of the execution process of LBS3, we select some examples from the GSM8K (Table 53 to Table LABEL:easy_hard_example_gsm8k_one_solve_target:), SQA (Table 58 to Table LABEL:easy_hard_example_sqa_two_solve_target:), and BBH-raco (Table 63 to Table LABEL:easy_hard_example_bbh_two_solve_target:) benchmarks in Section 4.3.2 for demonstration.
Instruction: Perform the difficulty rating task with the following steps:
1. **Read and Understand**: Carefully read both the original and target questions to ensure full comprehension.
2. **Compare**: Analyze and compare the two questions regarding their topics, complexity, required knowledge, and solution steps.
3. **Evaluate**: Assess the difficulty level of the target question relative to the original.
4. **Rate**:
- Assign 0 if both questions have similar difficulty.
- Assign -1 if the target question is slightly easier.
- Assign -2 if the target question is significantly easier.
- Assign 1 if the target question is slightly more difficult.
- Assign 2 if the target question is significantly more difficult.
Example:
Original Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?
Target Question: Jacob baked cookies for 36 of his neighbors in July, and then he baked twice as many cookies in August. How many cookies did Jacob bake altogether in July and August?
Rating: 0
**Provide the rating result without detailed explanation or analysis:**
[…]
Rating:
Table 3: Scoring rules of difficulty for proxy queries using GPT-4o
Appendix E LBS3 with Self-ICL and Auto-ICL
Appendix F Utility of Progressive Strategy
@@
Input:
Generate three different new analogous problems for the example problem below. These new problems are progressively more challenging to solve, and the most challenging new problem is no more difficult than the example problem below.
Example problem:
New problem 1:
Q:
Output:
Sam is observing a bird’s nest to see how many birds come and go in a day. He sees 20 birds leave the nest in the first 4 hours, and then he sees 1/2 that many birds return in the next 4 hours. He sees two times as many birds as he saw first leave the nest fly from the nest and leave in the next 4 hours. Then every bird that left before that hadn’t already returned returns to the nest in the next 4 hours. How many birds did Sam see return to the nest in the last 4 hours of the day?
New problem 2:
Q:
Linda is watching a squirrel’s nest to see how many squirrels come and go in a day. She sees 40 squirrels leave the nest in the first 8 hours, and then she sees 1/2 that many squirrels return in the next 8 hours. She sees two times as many squirrels as she saw first leave the nest fly from the nest and leave in the next 8 hours. Then every squirrel that left before that hadn’t already returned returns to the nest in the next 8 hours. How many squirrels did Linda see return to the nest in the last 8 hours of the day?
New problem 3:
Q:
John is observing a fish tank to see how many fish come and go from a particular corner in a day. He sees 50 fish leave the corner in the first 10 hours, and then he sees 1/2 that many fish return in the next 10 hours. He sees two times as many fish as he saw first leave the corner swim from the corner and leave in the next 10 hours. Then every fish that left before that hadn’t already returned returns to the corner in the next 10 hours. How many fish did John see return to the corner in the last 10 hours of the day?